A proton beam in an accelerator carries a current of 132
Answer #1

dQ= i dt where Q is charge i=current and t=time

we are given, that 132 uA flows in 1 second.

so, for 16 seconds, amount of charge = 132uA*16 = 2112uC = 2112 * 10^-6 C

Amount of charge on each proton = 1.6 * 10^-19 C

So, number of protons =(2112* 10^-6) / (1.6 * 10^-19 C) = 1320*10^13 . (say about 1.32*10^16 approx)
Answer #2

dQ= i dt where Q is charge i=current and t=time

we are given, that 132 uA flows in 1 second.

so, for 16 seconds, amount of charge = 132uA*16 = 2112uC = 2112 * 10^-6 C

Amount of charge on each proton = 1.6 * 10^-19 C

So, number of protons =(2112 * 10^-6) / (1.6 * 10^-19 C) = 1320*10^13
Answer #3

number of protons strike the target in a period of 16.0 s={(132*10-6)*16}/1.6*10-19=1320*1013=1.32*1016proton